How do you do surface area if you have a cylinder?
How do you do surface area if you have a rectangle?
Can someone show me some examples? THANKS A LOT!!!

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

How do you do surface area if you have a cylinder?
Formula: surface area = (circumference)*(height)
= 2(pi)r*h
---------------
How do you do surface area if you have a rectangle?
Formula: surface area = length*width
---------------
Post a couple of your problems and you will get help
using these formula.
===================
Cheers,
Stan H.

Answer by checkley71(8403)

SURFACE AREA OF A CYLINDER=C*H WHERE C=THE CIRCIMFERANCE OF THE BASE & H=HEIGHT.
SURFACE AREA OF A RECTANGLE=B1*B2*H WHERE B1=ONE OF THE BASE MEASUREMENTS, B2=
THE OTHER BASE MEASUREMENT & H=HEIGHT.